# A Probabilistic Parking Process and Labeled IDLA

### v1

@@ 2. Sampling Weakly Increasing Parking Functions and Lucky Cars
It is well known that the number of weakly increasing parking function of length $n$ is given by the $n$th Catalan number $C_{n}=\frac{1}{n+1}\binom{2n}{n}$. Therefore, weakly increasing parking functions are in bijection with the wide variety of Catalan objects [40]. For example, weakly increasing parking functions of length $n$ are in bijection with Dyck paths of length $2n$, as well as with binary trees with $n+1$ leaves. We recall that a Dyck path is a diagonal lattice path from the origin $(0,0)$ to the $x$-axis, consisting of only “up steps” $U=(1,1)$ and “down steps” $D=(1,-1)$. A Dyck path of length $2k\in\mathbb{N}$ involves $k$ up steps and $k$ down steps, and ends at $(0,2k)$. A classical bijection between Dyck paths of length $2n$ and weakly increasing parking functions of length $n$ sets the $i$th entry of the preference list to be $1$ plus the number of “down steps” occurring before the $i$th “up step.” We illustrate this bijective mapping in Figure 1.
@@
Therefore, to sample a uniformly random weakly increasing parking function, one can use of one of the many methods to generate uniformly random Catalan objects, such as Rèmy’s algorithm [37]. Refer to [16] for an algorithm to sample uniformly from the set of all parking functions.
@@ 2.1. Distribution of the Last Entry of a Weakly Increasing Parking Function
Let $\alpha\sim\mathcal{U}(\operatorname{PF}_{n})$ be sampled uniformly at random from the set of parking functions of length $n\in\mathbb{N}$. If $n$ is sufficiently large, then
$\Pr_{\alpha\,\sim\,\mathcal{U}(\operatorname{PF}_{n})}\left[\alpha\in\mathsf{PF}^{\uparrow}_{n}\right]\sim\frac{1}{e\sqrt{\pi n}}\left(\frac{4}{n}\right)^{n}.$
@@
Recall that $|\operatorname{PF}_{n}|=(n+1)^{n-1}$ and $|\mathsf{PF}^{\uparrow}_{n}|=\frac{1}{n+1}\binom{2n}{n}$. Therefore,
$\Pr_{\alpha\,\sim\,\mathcal{U}(\operatorname{PF}_{n})}\left[\alpha\in\mathsf{PF}^{\uparrow}_{n}\right]$ $=\frac{\frac{1}{n+1}\binom{2n}{n}}{(n+1)^{n-1}}=\frac{1}{(n+1)^{n}}\frac{(2n)!}{(n!)^{2}}\stackrel{{\scriptstyle(*)}}{{\sim}}\frac{1}{(n+1)^{n}}\frac{\sqrt{2\pi(2n)}\left(\frac{2n}{e}\right)^{2n}}{(2\pi n)\left(\frac{n}{e}\right)^{2n}}$
@@
where ()(*)( ∗ ) follows from Stirling’s approximation of the factorial. ∎
Diaconis and Hicks, in [16], study the distribution of the first entry of a uniformly random parking function. As motivated by our results stated in Section 1.2, an analogous study specifically for weakly increasing parking function is also of interest. However, in the weakly increasing setting, the first entry is always equal to $1$. Therefore, in what follows we focus on studying the distribution of the last entry $\alpha_{n}$ of a uniformly random weakly increasing parking function.
In order to compute distributional information, we establish the following result which has been simultaneously and independently derived by Butler et al. [3].
@@
  • If $\alpha\in\mathsf{PF}^{\uparrow}_{n}$, then $\alpha$ is in bijection with a Dyck path of length $2n$. Furthermore, if $\alpha_{i}=j$ then, in particular, the preference lists $(\alpha_{i+1},\alpha_{i+2},\ldots,\alpha_{n})$ are in bijection with North-East lattice paths from $(i+1,n-j+1)$ to $(n,n)$ that do not cross above the main diagonal. After a sequence of transformations one can obtain that these paths are in bijection with Dyck paths that start from $(0,0)$ and end at $(n-k+1,n-i)$. It is again well-known that these are counted by entries of the Catalan triangle [14], specifically by
@@
where ()(*)( ∗ ) and ()($\dagger$)( † ) follow from Stirling’s approximation of the factorial. ∎
@@
Let $\alpha\sim\mathcal{U}(\operatorname{PF}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ sufficiently large, and $j\geq 1$ fixed with respect to $n$,
$\Pr_{\alpha\sim\mathcal{U}(\operatorname{PF}_{n})}\left[\alpha_{n}=n-j\right]\sim\frac{1}{\sqrt{2}}\frac{j-1}{4^{n}}\,.$
@@
$\begin{split}\frac{(n+1)(j-1)}{n}\binom{2n-j-2}{n-j-1}\binom{2n}{n}^{-1}&=(j-1)\frac{(2n-j-2)!n!n!}{(n-j-1)!(n-1)(2n)!}\\ &\stackrel{{\scriptstyle(\ast)}}{{\sim}}{(j-1)}\frac{(2n-j-2)^{2n-j-2+\frac{1}{2}}\sqrt{2\pi}n^{n+\frac{1}{2}}\sqrt{2\pi}n^{n+\frac{1}{2}}\sqrt{2\pi}e^{4n+j+2}}{(n-j-1)^{n-j-1+\frac{1}{2}}\sqrt{2\pi}(n-1)^{n-1}\sqrt{2\pi}(2n)^{2n+\frac{1}{2}}\sqrt{2\pi}e^{4n+j+2}}\\ &\sim{(j-1)}\frac{(2n)^{2n-j-2+\frac{1}{2}}(1-\frac{j}{2n}-\frac{1}{n})^{2n-j-1+\frac{1}{2}}n^{n+\frac{1}{2}}n^{n+\frac{1}{2}}}{n^{n-j-1+\frac{1}{2}}(1-\frac{j}{n}-\frac{1}{n})^{n-j-1+\frac{1}{2}}n^{n-1+\frac{1}{2}}(1-\frac{1}{n})^{n-1+\frac{1}{2}}(2n)^{2n+\frac{1}{2}}}\\ &\sim{(j-1)}\frac{e^{-j-2}}{e^{-j-2}2^{2n+\frac{1}{2}}}\\ &\sim\frac{j-1}{4^{n}\sqrt{2}}\\ \end{split}$
where ()(*)( ∗ ) follows from Stirling’s approximation of the factorial. ∎
@@
Let $\alpha\sim\mathcal{U}(\operatorname{PF}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ and $j$ both going to infinity, we have
$\Pr_{\alpha\,\sim\,\mathcal{U}(\operatorname{PF}_{n})}\left[\alpha_{n}=j\right]\sim\frac{(n+1)e^{n+j-1}(n-j+1)\sqrt{2\pi n}}{(n+j-2)4^{n}}\Pr[X=j-1]\,.$
@@
$\stackrel{{\scriptstyle(\ast)}}{{\sim}}{(n-j+1)(n+1)}\frac{(n+j-2)^{n+j-2+\frac{1}{2}}\sqrt{2\pi}n^{n}\sqrt{2\pi n}e^{2n}}{\sqrt{2\pi}(j-1)!(2n)^{2n+\frac{1}{2}}e^{2n+j-2}}$
$\sim\frac{(n+1)(n-j+1)n^{n+\frac{1}{2}}(1+\frac{j-2}{n})^{n+\frac{1}{2}}(n+j-2)^{j-2}n^{n}\sqrt{2\pi n}e^{2n}}{(j-1)!4^{n+\frac{1}{2}}(n)^{2n+\frac{1}{2}}e^{2n+j-2}}$
$\sim\frac{(n+1)(n-j+1)e^{j-2}(n+j-2)^{j-2}\sqrt{2\pi n}}{(j-1)!4^{n+\frac{1}{2}}e^{j-2}}$
$=\frac{(n+1)e^{n+j-1}(n-j+1)\sqrt{2\pi n}}{(n+j-2)4^{n+\frac{1}{2}}}\Pr[X=j-1]$
where, once more, ()(*)( ∗ ) follows from Stirling’s approximation of the factorial. ∎
@@
$\begin{split}\mathbb{E}\left[\,\alpha_{n}\right]&={(n+1)}\binom{2n}{n}^{-1}\left[\sum\limits_{j=0}^{n-1}\frac{j(n-j)}{n}\binom{n+j-1}{j}\right]\\ &={(n+1)}\binom{2n}{n}^{-1}\left[\sum\limits_{j=0}^{n-1}j\binom{n+j-1}{j}-\frac{1}{n}\sum\limits_{j=0}^{n-1}j^{2}\binom{n+j-1}{j}\right].\end{split}$
@@ 2.2. Lucky Cars in Weakly Increasing Parking Functions
For completeness, in this subsection we bring together some recent enumerative results on parking function stating them in the language of probability. Harris and Martinez [27] previously considered the following enumerative question: Given a subset $L\subseteq[n]$ containing $1$, how many parking functions in $\operatorname{PF}_{n}$ have their set of lucky cars being those indexed by the set $L$? Their results involve enumerating these parking functions based on the outcome permutation, namely the permutation which details which car parked in what spot. They then restrict to the set of weakly increasing parking functions and ask the analogous question. Their result in this case involves a product of Catalan numbers, see [27, Theorem 2.8].
@@
Deutsch [14] showed that expected number of returns for a Dyck path of semilength $n$ (length $2n$) is $\frac{3n}{n+2}$. Thus, the expected number of lucky cars in a weakly increasing parking function is precisely $\frac{3n}{n+2}$. ∎
@@
The case of $1\not\in L$ is clear, as mentioned previously, assuming we run the parking protocol with a weakly increasing parking function, the first car is always lucky. When $1\in L$, the result follows from [27, Theorem 2.8] which states that the number of weakly increasing parking functions with lucky cars in the set $L$ is exactly
@@
where $C_{x_{j}}$ is the $x_{j}$th Catalan number and $x_{j}\coloneqq l_{j+1}-l_{j}-1$ and $x_{k}\coloneqq n-l_{k}$ when $1\in L$. This result also follows from the bijection to Dyck paths; the number of parking functions with lucky cars given by the set $L$ are in bijection with the set of Dyck paths with returns with the set of coordinates $(2l-2,0)$ for each $l\in L\setminus\{1\}$. We can count the number of such Dyck paths by thinking of each Dyck paths as a composition of smaller Dyck paths with exactly 1 return and length equal to the difference in position of the returns in our original Dyck path (see Figure 2). Our result follows from the fact that the number of Dyck paths with semilength $l_{j+1}-l_{j}$ and exactly 1 return is precisely $C_{l_{j+1}-l_{j}}$.
$x$$y$$(4,0)$$(10,0)$
Figure 2. The Dyck path corresponding to $(11334)\in\mathsf{PF}^{\uparrow}_{5}$ can be thought of as a composition of the two different Dyck paths, each with no returns except the last.
@@
It is well-known (see [14]) that the number of Dyck paths with $k$ returns is
@@ 3. Probabilistic Parking in the Unbounded Model
spots before parking in spot $i$. Now, consider the arrival of car $i$ under the unbounded stochastic model, assuming spots $11,2,\ldots,i-$ 1 are all occupied. In this case, under the unbounded model car $i$ (and thus $X_{\alpha}^{i}=1$) if and only if car $i$ hits position $i$ before drifting off to infinity on the line. As it turns out, the analysis for this single-car setting reduces to the analysis of McMullen [34] for the “bust time” of a gambler with some initial endowment. In particular, McMullen considers the following classical gambler’s ruin set up. A gambler with an initial endowment of $k$ dollars begins to place bets in a game where, with probability $p$, the gambler wins one dollar, and with probability $q=1-p$, they lose one dollar. McMullen is interested in evaluating the function $c(b,k)$ which is the number of ways the gambler can become ruined (i.e., be left with zero dollars) after starting to gamble with $k$ dollars. McMullen showed that $c(b,k)$ satisfies the recurrence relation
(3.1) $c(b,k)=c(b-1,k-1)+c(b-1,k+1)\,,$
@@
with boundary conditions $c(0,0)=1$, $c(\kappa,\kappa)=1$, and $c(\kappa,0)=0$. One can also think of this as the number of paths on $\mathbb{Z}^{+}$ of length $b$ starting from position $k$ which stay positive until ending at position 00.
@@
Translating this to our probabilistic parking process, when $\alpha$ is a parking function whose outcome is the identity permutation, car $i$ starts its random walk on $\mathbb{Z}$ in spot $\alpha_{i}\in[i]$, and it must displace $d_{i}$ spots to the right of spot $\alpha_{i}$ before parking in spot $i$. In the best case scenario, in total, car $i$ displaces to left exactly 00 times and to right exactly $d_{i}$ times, for a combined displacement of $d_{i}$. This can happen in exactly $1$ way. In the next best case scenario, in total, car $i$ displaces to the left exactly $1$ time and to the right exactly $d_{i}+1$ times, for a combined displacement of $d_{i}+2$. This can happen in exactly $d_{i}$ ways. In general, if car $i$ displaces to the left exactly $\ell\in\mathbb{N}$ times, then it must displace to the right exactly $d_{i}+\ell$ times for a combined displacement of $d_{i}+2\ell$.
@@
One can see that the number of ways in which car $i$ displaces a total of $\kappa\geq 0$ times, starting from $\delta\leq\kappa$ spots to the left of spot $i$, before parking in spot $i$ is exactly characterized by the recurrence relation given in (3.1). We denote this number by $c^{i}(\kappa,\delta)$. In particular, we have that car $i$ parks in spot $i$ after taking $d_{i}+2\ell$ steps of a random walk starting from position $\alpha_{i}$ if and only if a gambler becomes ruined after $d_{i}+2\ell$ rounds of a game starting with $d_{i}$ dollars. Additionally, the probability of winning a dollar corresponds to the probability of a car taking a step to the left, which is given by $q$. Thus we are interested in the value of $c^{i}(d_{i}+2\ell,d_{i})$ with the same boundary conditions as in the gambler’s ruin scenario. McMullen [34] showed that that the recurrence relation (3.1) unravels to obtain
@@
which is the $\ell$th term of the $d_{i}$th Catalan self-convolution [5]. From this, we obtain the following.
@@
In Corollary 3.3, several of the results make use of the total displacement statistic (the sum of the deterministic displacement of each of the cars) showing that it is indeed central to determining the probability, expected time, and variance of the probabilistic parking process. For an overview of results related to the total displacement statistic and parking functions, see [44].

### v2

@@ 2. Sampling Weakly Increasing Parking Functions and Lucky Cars
It is well known that the number of weakly increasing parking function of length $n$ is given by the $n$th Catalan number $C_{n}=\frac{1}{n+1}\binom{2n}{n}$. Therefore, weakly increasing parking functions are in bijection with the wide variety of Catalan objects [stanley2015catalan]. For example, weakly increasing parking functions of length $n$ are in bijection with Dyck paths of length $2n$, as well as with binary trees with $n+1$ leaves. We recall that a Dyck path is a diagonal lattice path from the origin $(0,0)$ to the $x$-axis, consisting of only “up steps” $U=(1,1)$ and “down steps” D=(1,1)D=(1,-1). A Dyck path of length $2k\in\mathbb{N}$ involves $k$ up steps and $k$ down steps, and ends at $(0,2k)$. A classical bijection between Dyck paths of length $2n$ and weakly increasing parking functions of length $n$ sets the $i$th entry of the preference list to be $1$ plus the number of “down steps” occurring before the $i$th “up step.” We illustrate this bijective mapping in Figure 2.
@@
Therefore, to sample a uniformly random weakly increasing parking function, one can use of one of the many methods to generate uniformly random Catalan objects, such as Rèmy’s algorithm [Remy]. Refer to [diaconis2017probabilizing] for an algorithm to sample uniformly from the set of all parking functions.
@@ 2.1. Distribution of the Last Entry of a Weakly Increasing Parking Function
Let $\alpha\sim\mathcal{U}(\mathsf{PF}_{n})$ be sampled uniformly at random from the set of parking functions of length $n\in\mathbb{N}$. If $n$ is sufficiently large, then
$\Pr_{\alpha\,\sim\,\mathcal{U}(\mathsf{PF}_{n})}\left[\alpha\in\mathsf{PF}^{\uparrow}_{n}\right]\sim\frac{1}{e\sqrt{\pi n}}\left(\frac{4}{n}\right)^{n}.$
@@
Recall that $|\mathsf{PF}_{n}|=(n+1)^{n-1}$ and $|\mathsf{PF}^{\uparrow}_{n}|=\frac{1}{n+1}\binom{2n}{n}$. Therefore,
$\Pr_{\alpha\,\sim\,\mathcal{U}(\mathsf{PF}_{n})}\left[\alpha\in\mathsf{PF}^{\uparrow}_{n}\right]$ $=\frac{\frac{1}{n+1}\binom{2n}{n}}{(n+1)^{n-1}}=\frac{1}{(n+1)^{n}}\frac{(2n)!}{(n!)^{2}}\stackrel{{\scriptstyle(*)}}{{\sim}}\frac{1}{(n+1)^{n}}\frac{\sqrt{2\pi(2n)}\left(\frac{2n}{e}\right)^{2n}}{(2\pi n)\left(\frac{n}{e}\right)^{2n}}$
@@
where ()(*) follows from Stirling’s approximation of the factorial. ∎
Diaconis and Hicks, in [diaconis2017probabilizing], study the distribution of the first entry of a uniformly random parking function. As motivated by our results stated in Section 1.2, an analogous study specifically for weakly increasing parking function is also of interest. However, in the weakly increasing setting, the first entry is always equal to $1$. Therefore, in what follows we focus on studying the distribution of the last entry $\alpha_{n}$ of a uniformly random weakly increasing parking function.
In order to compute distributional information, we establish the following result which has been simultaneously and independently derived by Butler et al. [butler2024lucky].
@@
  • If $\alpha\in\mathsf{PF}^{\uparrow}_{n}$, then $\alpha$ is in bijection with a Dyck path of length $2n$. Furthermore, if $\alpha_{i}=j$ then, in particular, the preference lists $(\alpha_{i+1},\alpha_{i+2},\ldots,\alpha_{n})$ are in bijection with North-East lattice paths from (i+1,nj+1)(i+1,n-j+1) to $(n,n)$ that do not cross above the main diagonal. After a sequence of transformations one can obtain that these paths are in bijection with Dyck paths that start from $(0,0)$ and end at (nk+1,ni)(n-k+1,n-i). It is again well-known that these are counted by entries of the Catalan triangle [DyckPathEnumeration1999], specifically by
@@
where ()(*) and $(\dagger)$ follow from Stirling’s approximation of the factorial. ∎
@@
Let $\alpha\sim\mathcal{U}(\mathsf{PF}^{\uparrow}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ sufficiently large, and $j\geq 1$ fixed with respect to $n$,
$\Pr_{\alpha\sim\mathcal{U}(\mathsf{PF}_{n})}\left[\alpha_{n}=n-j\right]\sim\frac{1}{\sqrt{2}}\frac{j-1}{4^{n}}\,.$
@@
[…]
where ()(*) follows from Stirling’s approximation of the factorial. ∎
@@
Let $\alpha\sim\mathcal{U}(\mathsf{PF}^{\uparrow}_{n})$ be sampled uniformly at random from the set of all weakly increasing parking functions of length $n\in\mathbb{N}$. For $n$ and $j$ both going to infinity, we have
$\Pr_{\alpha\,\sim\,\mathcal{U}(\mathsf{PF}_{n})}\left[\alpha_{n}=j\right]\sim\frac{(n+1)e^{n+j-1}(n-j+1)\sqrt{2\pi n}}{(n+j-2)4^{n}}\Pr[X=j-1]\,.$
@@
$\hskip-90.3375pt\stackrel{{\scriptstyle(\ast)}}{{\sim}}{(n-j+1)(n+1)}\frac{(n+j-2)^{n+j-2+\frac{1}{2}}\sqrt{2\pi}n^{n}\sqrt{2\pi n}e^{2n}}{\sqrt{2\pi}(j-1)!(2n)^{2n+\frac{1}{2}}e^{2n+j-2}}$
$\hskip-90.3375pt\sim\frac{(n+1)(n-j+1)n^{n+\frac{1}{2}}(1+\frac{j-2}{n})^{n+\frac{1}{2}}(n+j-2)^{j-2}n^{n}\sqrt{2\pi n}e^{2n}}{(j-1)!4^{n+\frac{1}{2}}(n)^{2n+\frac{1}{2}}e^{2n+j-2}}$
$\hskip-90.3375pt\sim\frac{(n+1)(n-j+1)e^{j-2}(n+j-2)^{j-2}\sqrt{2\pi n}}{(j-1)!4^{n+\frac{1}{2}}e^{j-2}}$
$\hskip-90.3375pt=\frac{(n+1)e^{n+j-1}(n-j+1)\sqrt{2\pi n}}{(n+j-2)4^{n+\frac{1}{2}}}\Pr[X=j-1]$
where, once more, ()(*) follows from Stirling’s approximation of the factorial. ∎
@@
[…]
@@ 2.2. Lucky Cars in Weakly Increasing Parking Functions
For completeness, in this subsection we bring together some recent enumerative results on parking function stating them in the language of probability. Harris and Martinez [harris2024parkingfunctionsfixedset] previously considered the following enumerative question: Given a subset $L\subseteq[n]$ containing $1$, how many parking functions in $\mathsf{PF}_{n}$ have their set of lucky cars being those indexed by the set $L$? Their results involve enumerating these parking functions based on the outcome permutation, namely the permutation which details which car parked in what spot. They then restrict to the set of weakly increasing parking functions and ask the analogous question. Their result in this case involves a product of Catalan numbers, see [harris2024parkingfunctionsfixedset, Theorem 2.8].
@@
Deutsch [DyckPathEnumeration1999] showed that expected number of returns for a Dyck path of semilength $n$ (length $2n$) is $\frac{3n}{n+2}$. Thus, the expected number of lucky cars in a weakly increasing parking function is precisely $\frac{3n}{n+2}$. ∎
@@
The case of $1\not\in L$ is clear, as mentioned previously, assuming we run the parking protocol with a weakly increasing parking function, the first car is always lucky. When $1\in L$, the result follows from [harris2024parkingfunctionsfixedset, Theorem 2.8] which states that the number of weakly increasing parking functions with lucky cars in the set $L$ is exactly
@@
where $C_{x_{j}}$ is the $x_{j}$th Catalan number and $x_{j}\coloneqq l_{j+1}-l_{j}-1$ and $x_{k}\coloneqq n-l_{k}$ when $1\in L$. This result also follows from the bijection to Dyck paths; the number of parking functions with lucky cars given by the set $L$ are in bijection with the set of Dyck paths with returns with the set of coordinates (2l2,0)(2l-2,0) for each $l\in L\setminus\{1\}$. We can count the number of such Dyck paths by thinking of each Dyck paths as a composition of smaller Dyck paths with exactly 1 return and length equal to the difference in position of the returns in our original Dyck path (see Figure 3). Our result follows from the fact that the number of Dyck paths with semilength $l_{j+1}-l_{j}$ and exactly 1 return is precisely $C_{l_{j+1}-l_{j}}$.
xxyy(4,0)(4,0)$(10,0)$
Figure 3. The Dyck path corresponding to $(1,1,3,3,4)\in\mathsf{PF}^{\uparrow}_{5}$ can be thought of as a composition of the two different Dyck paths, each with no returns except the last.
@@
It is well-known (see [DyckPathEnumeration1999]) that the number of Dyck paths with $k$ returns is
@@ 3. Probabilistic Parking in the Unbounded Model
spots before parking in spot $i$. Now, consider the arrival of car $i$ under the unbounded stochastic model, assuming spots $1,2,\ldots,i-1$ are all occupied. In this case, under the unbounded model car $i$ (and thus $X_{\alpha}^{i}=1$) if and only if car $i$ hits position $i$ before drifting off to infinity on the line. As it turns out, the analysis for this single-car setting reduces to the analysis of McMullen [mcmullen2020drunken] for the “bust time” of a gambler with some initial endowment. In particular, McMullen considers the following classical gambler’s ruin set up. A gambler with an initial endowment of $k$ dollars begins to place bets in a game where, with probability $p$, the gambler wins one dollar, and with probability q=1pq=1-p, they lose one dollar. McMullen is interested in evaluating the function $c(b,k)$ which is the number of ways the gambler can become ruined (i.e., be left with zero dollars) after starting to gamble with $k$ dollars. McMullen showed that $c(b,k)$ satisfies the recurrence relation
(3.1) c(b,k)=c(b1,k1)+c(b1,k+1),c(b,k)=c(b-1,k-1)+c(b-1,k+1)\,,
@@
with boundary conditions $c(0,0)=1$, $c(\kappa,\kappa)=1$, and $c(\kappa,0)=0$. One can also think of this as the number of paths on $\mathbb{Z}^{+}$ of length $b$ starting from position $k$ which stay positive until ending at position 0.
@@
Translating this to our probabilistic parking process, when $\alpha$ is a parking function whose outcome is the identity permutation, car $i$ starts its random walk on $\mathbb{Z}$ in spot $\alpha_{i}\in[i]$, and it must displace $d_{i}$ spots to the right of spot $\alpha_{i}$ before parking in spot $i$. In the best case scenario, in total, car $i$ displaces to left exactly 0 times and to right exactly $d_{i}$ times, for a combined displacement of $d_{i}$. This can happen in exactly $1$ way. In the next best case scenario, in total, car $i$ displaces to the left exactly $1$ time and to the right exactly $d_{i}+1$ times, for a combined displacement of $d_{i}+2$. This can happen in exactly $d_{i}$ ways. In general, if car $i$ displaces to the left exactly $\ell\in\mathbb{N}$ times, then it must displace to the right exactly $d_{i}+\ell$ times for a combined displacement of $d_{i}+2\ell$.
@@
One can see that the number of ways in which car $i$ displaces a total of $\kappa\geq 0$ times, starting from $\delta\leq\kappa$ spots to the left of spot $i$, before parking in spot $i$ is exactly characterized by the recurrence relation given in (3.1). We denote this number by $c^{i}(\kappa,\delta)$. In particular, we have that car $i$ parks in spot $i$ after taking $d_{i}+2\ell$ steps of a random walk starting from position $\alpha_{i}$ if and only if a gambler becomes ruined after $d_{i}+2\ell$ rounds of a game starting with $d_{i}$ dollars. Additionally, the probability of winning a dollar corresponds to the probability of a car taking a step to the left, which is given by $q$. Thus we are interested in the value of $c^{i}(d_{i}+2\ell,d_{i})$ with the same boundary conditions as in the gambler’s ruin scenario. McMullen [mcmullen2020drunken] showed that that the recurrence relation (3.1) unravels to obtain
@@
[…]
@@
In Corollary 3.3, several of the results make use of the total displacement statistic (the sum of the deterministic displacement of each of the cars) showing that it is indeed central to determining the probability, expected time, and variance of the probabilistic parking process. For an overview of results related to the total displacement statistic and parking functions, see [yan2015parking].